OpenStudy (anonymous):

how do i use the geometric sequence formula to figure out what the 6th term is (after 11)? 15 14 13 12 11

OpenStudy (anonymous):

\[a_n=a_1r ^{n-1} \] this formula

OpenStudy (anonymous):

First you need to find r, the common ratio do you know how to find this?

OpenStudy (anonymous):

is this geometric

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would i find the ratio by dividing each one?

OpenStudy (anonymous):

14/15=.93333333 13/14=.92857.... you should get the same number if it's geometric. You divide a term by the previous one

OpenStudy (anonymous):

ok, so do i round it off to .93 for all of them? what do i do then?

OpenStudy (anonymous):

I'm not sure this is geometric but if it were you would simply use your formula that says that for any term a_n, in this case the 6th term is when n=6, you multiply the first term a1 by the common ratio r to the power of n-1=6-1=5 So a_6=15*(r)^5

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

i did 15*9.3^5 and got this weird number. 1043532.554

OpenStudy (anonymous):

you get a number like 10.435.... because you have r=.93 not 9.3 i still don't think you have a geometric here you seem to have an arithmetic sequence so your next term is 10.

OpenStudy (anonymous):

it´s supposed to be a geometric sequence.

OpenStudy (anonymous):

How about for this geometric sequence: 15,13,11,9,7 ? i got a weird answer

OpenStudy (anonymous):

@marybel86